# Parsimonious Gaussian mixture models with piecewise-constant eigenvalue profiles

### v1

@@ 4.1 Componentwise monotonicity implies mixture monotonicity
It is well known from the classical EM theory [27] that increasing the expected complete-data log-likelihood $\mathbb{E}_{\,y\mid x,\theta\,}[\ln\mathcal{L}_{\mathrm{c}}(\cdot)]$ (16) increases the mixture log-likelihood $\ln\mathcal{L}(\cdot)$ (14). This is notably due to Gibbs’ inequality—i.e. the non-negativity of the relative entropy. In this work, we are not interested in increasing the log-likelihood but a penalized version of it (23), to enforce parsimonious modeling. As said earlier, we would like to select the eigenvalues multiplicities with componentwise model selection strategies drawn from the PSA methodology. However, such a variant of the EM algorithm would not enjoy a priori any guarantee on the evolution of the objective function (23) over the iterations. The following theorem states that a thoroughly-designed variant of the EM algorithm where we replace the M-step with a componentwise model selection guarantees the monotonicity of the penalized log-likelihood. We call this variant componentwise penalized EM (CPEM).
@@
Remark 3.
Using Theorem 2, one can show with a similar reasoning as in Equation (6) that ${}{\Psi_{c}(\gamma_{c})=-\frac{n\pi_{c}}{2}\sum_{k=1}^{d_{c}}\gamma_{ck}\ln\hat{\lambda}_{ck}(t_{c},\gamma_{c})+\mathrm{cst}}$, which simplifies the implementation. For instance, with a BIC-like penalization ($\alpha=\ln(n)/2$), the problem of model selection (24) can simply be formulated as
@@
Remark 4.
Theorem 4 applies straightforwardly to other parsimonious GMMs such as the mixtures of PPCA of [61], the HDDC models of [11] and the parsimonious models of [48] (which the MPSA models generalize), by choosing types of the form $\gamma_{c}=(1,\dots,1,p-q_{c})$. Therefore, our results implicitly address the important issue of automatic intrinsic dimension selection in low-rank mixture models that we raised in Section 1. At each step of the EM, for each mixture component $c$, we could select the intrinsic dimension $q_{c}$ that solves
@@
Remark 5.
As far as we know, our componentwise penalized EM monotonicity theorem is new. We believe that it can have an important impact in the parsimonious mixture models community [61, 48, 57]. Indeed, the problem of hyperparameter selection in mixture models is notoriously challenging and open, and it is often addressed heuristically without theoretical guarantees [11]. Therefore, the monotonicity result that we obtain addresses this important issue. Let us note that the idea of integrating both parameter estimation and model selection in an EM algorithm is absolutely not new; it was notably proposed in the seminal work of [29] for finite mixture models and theoretically investigated in [41] for missing data. Let us also note that our algorithm has little to do with the componentwise EM of [23], which updates the mixture parameters sequentially, one component per iteration.
@@ 4.2 Algorithm
Theorem 4 motivates us to design a variant of the EM algorithm where the componentwise quantities $\Psi_{c}(\cdot)-\alpha\,\kappa(\cdot)$ increase at each EM iteration. For that, at each step $s$, for each mixture component $c$, we consider a set of candidate models $\mathcal{M}(\gamma_{c})\in\mathcal{P}(\mathcal{C}(p))$ and take the one that maximizes $\Psi_{c}(\cdot)-\alpha\,\kappa(\cdot)$. The general methodology is summarized in Algorithm 2.
@@
$\hat{\gamma}\in\mathcal{C}(p)^{C},\hat{\theta}\in\Theta(\hat{\gamma})$
$y\leftarrow\operatorname{KMeans}(x,C)\in{[\![1,C]\!]}^{n}$triangleright K-Means init.
$t\leftarrow\left(\mathbf{1}_{y_{i}=c}\right)_{c,i}$triangleright posterior proba.
$\hat{\gamma}\leftarrow\gamma_{0}$triangleright types init. (spherical or full)
@@
     $\hat{\gamma}\leftarrow\operatorname*{arg\,max}\limits_{\gamma_{c}\in\{\hat{\gamma}_{c}\}\cup\mathcal{M}(\hat{\gamma}_{c})}\Psi_{c}(\gamma_{c})-\alpha\,\kappa(\gamma_{c})$triangleright thm. 4
     $\hat{\theta}\leftarrow\operatorname{M-step}(t,\hat{\gamma})$triangleright thm. 3
     $t\leftarrow\operatorname{E-step}(\hat{\theta})$triangleright thm. 2
@@
Remark 6.
Note that at each step, we include the current model $\hat{\gamma}_{c}$ into the set of candidate models in order to make sure that there is at least one model that increases the penalized log-likelihood. Indeed, Equation (24) is always satisfied for $\gamma^{\prime}_{c}=\gamma_{c}$.
@@
[…]
Remark 7.
Early experiments showed that the above-defined bottom-up and top-down strategies may get stuck in local optima due to their monotonous behavior in terms of model complexity. For instance, the bottom-up approach may increase too much the number of parameters during the first few iterations (similarly as in [36], Figure 7) because one needs a lot of degrees of freedom to model the density when the mixture centers are initialized far from the true cluster centers. Once the true centers are found, one may realize that the clusters have a much simpler structure and one would like to reduce the number of parameters (as we can see on the bottom of Figure 2 for the red curve). To enable such late changes, we eventually take $\mathcal{M}(\gamma_{c})=\mathcal{N}^{+}(\gamma_{c})\cup\mathcal{N}^{-}(\gamma_{c})$ for both bottom-up and top-down strategies.
@@ 5 Experiments
In this section, we perform various density modeling and clustering experiments with our new MPSA models (MPSA-H for the hierarchical clustering strategy, MPSA-R for the relative eigengap strategy, MPSA-U for the bottom-up strategy and MPSA-D for the top-down strategy), and compare them to the more standard full Gaussian mixture model (GMM-F) and the spherical one (GMM-S).
@@ 5.1 CPEM in action for 2D data
[…]
@@ 5.2 Model selection strategies
Figure 2: Evolution of the penalized-likelihood and the number of parameters over the iterations of the CPEM algorithm for the four model selection strategies (MPSA-H/R/U/D). Each transparent curve refers to one repetition while the thick opaque curve represents the average over the 10 independent repetitions.
We see that the penalized-likelihoods are once again monotonous, as expected by theory (cf. Theorem 4).
Let us now comment the results based on our intuitions. All penalized-likelihood curves seem to converge to the same optimum, after approximately 5 to 20 iterations. MPSA-D takes a much larger number of iterations to converge than the other strategies. However, the learning curves (penalized-likelihood and number of parameters) are quite regular and with small variance, compared to the other strategies. We guess that in higher dimensions, MPSA-D may badly behave at first iterations due to the initialization with full covariance matrices, and it can moreover require a large number of iterations to converge, since it can only decrease the model complexity one by one. Therefore, smarter initialization heuristics could be designed, but this is left to future research. MPSA-U also yields quite regular learning curves, although we can observe small oscillations of the number of parameters near the optimum (cf. Remark 7 in Section 4.2). We moreover guess that MPSA-U may get stuck in local optima due to the simplicity of the models at first iterations [65]. MPSA-H and MPSA-R are more flexible in terms of model complexity—which may make them attain the optimum faster as we can see in Figure 2. However, we guess that their flexibility may also cause higher variance in performance and parsimony.
5.3 Density modeling benchmark
The goal of this experiment is to compare our new parsimonious GMMs (MPSA-H/R/U/D) to the more-standard GMMs (full and spherical) in terms of penalized log-likelihood (23) $\sum_{i=1}^{n}\ln({\sum_{c=1}^{C}\hat{\pi}_{c}\,\mathcal{N}(\,x_{i}\mid\hat{\mu}_{c},\hat{\Sigma}_{c}\,)})-\alpha\,\kappa(\hat{\gamma})$ (with $\alpha=\ln(n)/2$) under various settings. All datasets but the last one are synthetic, sampled according to the generative model presented in the beginning of this section. The hyperparameters are $C=3$, $n=1000$, $(\pi_{1},\pi_{2},\pi_{3})=(0.4,0.3,0.3)$, $\mu_{c}\sim\mathcal{U}\left(\left[-5,5\right]^{p}\right)$, $(\lambda_{11},\lambda_{21},\lambda_{31})=(3,2,1)$, $\lambda_{cp}/\lambda_{c1}=0.01\,\forall c$, and $p\in\{10,100\}$. For the first two datasets (MPSA $10$ and MPSA $100$), one takes ${\gamma_{1},\gamma_{2},\gamma_{3}=(1,p-1),(1,2,p-3),(1,2,4,p-7)}$ for $p\in\{10,100\}$. For the third and fourth dataset (Full $10$ and Full $100$), one takes ${(\gamma_{1},\gamma_{2},\gamma_{3})=((1^{p}),(1^{p}),(1^{p}))}$ for $p\in\{10,100\}$, i.e. full Gaussian models with distinct population eigenvalues. The Skew $100$ dataset has ${(\gamma_{1},\gamma_{2},\gamma_{3})=((1^{p}),(1^{p}),(1^{p}))}$ for $p=100$ and a skew-normal distribution (instead of the Gaussian one used in the previous experiments). The last dataset (Ionosphere) is real [60], with $C=2$, $n=351$ and $p=34$. We report in Table 1 the average penalized-likelihood reached at the end of the CPEM algorithm, over 10 independent $n$-samples of the synthetic datasets. The last dataset—which is real—is split using a stratified $10$-fold cross-validation.
Table 1: Comparison of the penalized-likelihoods $\sum_{i=1}^{n}\ln({\sum_{c=1}^{C}\hat{\pi}_{c}\,\mathcal{N}(\,x_{i}\mid\hat{\mu}_{c},\hat{\Sigma}_{c}\,)})-\alpha\,\kappa(\hat{\gamma})$ reached by the MPSA-H/R/U/D models at the end of the CPEM algorithm, and the standard GMMs under various data distributions. The leading penalized-likelihoods are in bold.
Model MPSA $10$ MPSA $50$ Full $10$ Full $100$ Skew $100$ Ionosphere
@@ 5.3 Density modeling benchmark
MPSA-R $\mathbf{-0.65\pm 0.06}$ $\mathbf{46.00\pm 0.22}$ $\mathbf{-07.59\pm 0.06}$ $-100.02\pm 0.28$ $-099.09\pm 0.72$ $\mathbf{+04.76\pm 1.67}$
@@
In the first setting (MPSA $10$), all our MPSA models significantly outperform the classical GMMs. Since $p=10$ and $n=1000$, one probably has enough samples to estimate the full (anisotropic) covariance parameters, and the additional parsimony of our MPSA models yields a significantly higher penalized-likelihood. In the second setting (MPSA $100$), all our MPSA models again significantly outperform the classical GMMs. Since $p=100$ and $n=1000$, one probably does not have anymore enough samples to estimate the full covariance parameters, therefore the full GMM begins to severely drop in performance, while the spherical model is not flexible enough to model the data distribution due to its isotropic nature. In the third setting (Full $10$), since the true model is full and we have enough samples, the full GMM is the best estimator, and our MPSA models are on par with it, while the spherical GMM is below. In the fourth setting (Full $100$), the true model is still full but we don’t have enough samples anymore, so that the full GMM gets worse and the spherical GMM becomes better. Except for MPSA-R which is on par with GMM-S, our other MPSA models are significantly better. The same kind of observation goes for the fifth setting (Skew $100$). In the sixth setting (Ionosphere), our MPSA models significantly outperform the other GMMs.

### v2

@@ 4.1 Componentwise monotonicity implies mixture monotonicity
It is well known from the classical EM theory [dempster_maximum_1977] that increasing the expected complete-data log-likelihood $\mathbb{E}_{\,y\mid x,\theta\,}[\ln\mathcal{L}_{\mathrm{c}}(\cdot)]$ (16) increases the mixture log-likelihood $\ln\mathcal{L}(\cdot)$ (14). This is notably due to Gibbs’ inequality—i.e. the non-negativity of the relative entropy. In this work, we are not interested in increasing the log-likelihood but a penalized version of it (23), to enforce parsimonious modeling. As said earlier, we would like to select the eigenvalues multiplicities with componentwise model selection strategies drawn from the PSA methodology. However, such a variant of the EM algorithm would not enjoy a priori any guarantee on the evolution of the objective function (23) over the iterations. The following theorem states that a thoroughly-designed variant of the EM algorithm where we replace the M-step with a componentwise model selection guarantees the monotonicity of the penalized log-likelihood. We call this variant componentwise penalized EM (CPEM).
@@
Remark 3 (Simplification of componentwise inequalities).
Using Theorem 2, one can show with a similar reasoning to Equation (6) that $\Psi_{c}(\gamma_{c})=-\frac{n\pi_{c}}{2}\sum_{k=1}^{d_{c}}\gamma_{ck}\ln\hat{\lambda}_{ck}(t_{c},\gamma_{c})+\mathrm{cst}$}, which simplifies the implementation. For instance, with a BIC-like penalization ($\alpha=\ln(n)/2$), the problem of model selection (24) can simply be formulated as
@@
Remark 4 (Impact on low-rank mixture models).
Theorem 4 applies straightforwardly to other parsimonious GMMs such as the mixtures of PPCA of [tipping_mixtures_1999], the HDDC models of [bouveyron_high-dimensional_2007] and the parsimonious models of [mcnicholas_parsimonious_2008] (which the MPSA models generalize), by choosing types of the form $\gamma_{c}=(1,\dots,1,p-q_{c})$. Therefore, our results implicitly address the important issue of automatic intrinsic dimension selection in low-rank mixture models that we raised in Section 1. At each step of the EM, for each mixture component $c$, we could select the intrinsic dimension $q_{c}$ that solves
@@
Remark 5 (Novelty of CPEM).
As far as we know, our componentwise penalized EM monotonicity theorem is new. We believe that it can have an important impact in the parsimonious mixture models community [tipping_mixtures_1999, mcnicholas_parsimonious_2008, scrucca_mclust_2016]. Indeed, the problem of hyperparameter selection in mixture models is notoriously challenging and open, and it is often addressed heuristically without theoretical guarantees [bouveyron_high-dimensional_2007]. Therefore, the monotonicity result that we obtain addresses this important issue. Let us note that the idea of integrating both parameter estimation and model selection in an EM algorithm is absolutely not new; it was notably proposed in the seminal work of [figueiredo_unsupervised_2002] for finite mixture models and theoretically investigated in [jiang_e-ms_2015] for missing data. Let us also note that our algorithm has little to do with the componentwise EM of [celeux_componentwise_2001], which updates the mixture parameters sequentially, one component per iteration.
@@ 4.2 Algorithm
Theorem 4 motivates us to design a variant of the EM algorithm where the componentwise quantities $\Psi_{c}(\cdot)-\alpha\,\kappa(\cdot)$ increase at each EM iteration. For that, at each step $s$, for each mixture component $c$, we consider a set of candidate models $\mathcal{M}(\gamma_{c})\subset\mathcal{C}(p)$ and take the one that maximizes $\Psi_{c}(\cdot)-\alpha\,\kappa(\cdot)$. The general methodology is summarized in Algorithm 2.
@@
[…]
$y\leftarrow\operatorname{KMeans}(x,C)\in{[\![1,C]\!]}^{n}$ $\triangleright$ K-Means init.
$t\leftarrow\left(\mathbf{1}_{y_{i}=c}\right)_{c,i}$ $\triangleright$ posterior proba.
$\hat{\gamma}\leftarrow\gamma_{0}$ $\triangleright$ types init. (spherical or full)
@@
  $\hat{\gamma}\leftarrow\operatorname*{arg\,max}\limits_{\gamma_{c}\in\{\hat{\gamma}_{c}\}\cup\mathcal{M}(\hat{\gamma}_{c})}\Psi_{c}(\gamma_{c})-\alpha\,\kappa(\gamma_{c})$ $\triangleright$ thm. 4
[…]
@@
Remark 6 (Satisfying monotonicity conditions).
In order to ensure that the theoretical monotonicity conditions from Theorem 4 are met at each iteration—regardless of the set of candidate models $\mathcal{M}(\hat{\gamma}_{c})$—we include the current model $\hat{\gamma}_{c}$ in the set of candidate models. Indeed, not switching model at a CPEM iteration corresponds to a classical M-step from the EM algorithm and it naturally satisfies Equation (24) (with $\gamma^{\prime}_{c}=\gamma_{c}$). Consequently, the selected model, $\operatorname*{arg\,max}_{\gamma_{c}\in\{\hat{\gamma}_{c}\}\cup\mathcal{M}(\hat{\gamma}_{c})}\Psi_{c}(\gamma_{c})-\alpha\,\kappa(\gamma_{c})$ satisfies the theoretical monotonicity conditions too, a fortiori.
@@
We now introduce two additional ideas, borrowing from classical paradigms of model selection and sometimes referred to as top-down and bottom-up approaches (see [bach_optimization_2012] for uses in sparse optimization problems and [verbockhaven_growing_2024] for uses in neural architecture search). Let $\gamma\coloneqq(\gamma_{1},\dots,\gamma_{d})\in\mathcal{C}(p)$. We say that $\gamma^{\prime}\in\mathcal{C}(p)$ is an upper neighbor of $\gamma$ and write $\gamma^{\prime}\in\mathcal{N}^{+}(\gamma)$ if it can be expressed as $\gamma^{\prime}=(\gamma_{1},\dots,\gamma_{k-1},\gamma_{k1},\gamma_{k2},\gamma_{k+1},\dots,\gamma_{d})$ with $(\gamma_{k1},\gamma_{k2})\in\mathcal{C}(\gamma_{k})$, for any $k\in\{1,\dots,d\}$. Similarly, we say that $\gamma^{\prime}\in\mathcal{C}(p)$ is a lower neighbor of $\gamma$ and write $\gamma^{\prime}\in\mathcal{N}^{-}(\gamma)$ if it can be expressed as $\gamma^{\prime}=(\gamma_{1},\dots,\gamma_{k-1},\gamma_{k}+\gamma_{k+1},\gamma_{k+2},\dots,\gamma_{d})$. One can easily show that $\mathcal{N}^{+}(\gamma)=p-d$ and $\mathcal{N}^{-}(\gamma)=d-1$. [^4] These definitions are connected to the notion of type refinement introduced in [szwagier_curse_2025], and they can be more-intuitively visualized via the Hasse diagram representation of PSA models from the paper. The bottom-up strategy consists in starting from spherical GMMs ($\gamma_{c}=(p)\,\forall c$) and increasing step by step the model complexity by finding the “split” in the eigenvalue profile that maximizes the penalized log-likelihood (i.e. taking $\mathcal{M}(\gamma_{c})=\mathcal{N}^{+}(\gamma_{c})$). The top-down strategy is similar but instead starts from the full GMMs ($\gamma_{c}=(1,\dots,1)\,\forall c$) and decreases the complexity at each step by equalizing the two adjacent groups of eigenvalues that maximize the penalized log-likelihood (i.e. taking $\mathcal{M}(\gamma_{c})=\mathcal{N}^{-}(\gamma_{c})$).
Remark 7 (Up and down).
Early experiments showed that the above-defined bottom-up and top-down strategies may get stuck in local optima due to their monotonous behavior in terms of model complexity. For instance, the bottom-up approach may increase too much the number of parameters during the first few iterations (similarly to [houdard_high-dimensional_2018], Figure 7) because one needs a lot of degrees of freedom to model the density when the mixture centers are initialized far from the true cluster centers. Once the true centers are found, one may realize that the clusters have a much simpler structure and one would like to reduce the number of parameters (as we can see on the bottom of Figure 2 for the red curve). To enable such late changes, we eventually take $\mathcal{M}(\gamma_{c})=\mathcal{N}^{+}(\gamma_{c})\cup\mathcal{N}^{-}(\gamma_{c})$ for both bottom-up and top-down strategies.
Remark 8 (Which heuristic to choose?).
Four strategies for updating the eigenvalue multiplicities during the CPEM algorithm have been proposed in this subsection, all of which satisfy the theoretical monotonicity conditions (cf. Remark 6). While theory shows that each strategy yields a non-decreasing sequence of penalized log-likelihoods, it does not show how much the final parameters increase the penalized log-likelihood—with respect to initialization, to more classical GMMs, or in between the different strategies. Gibbs’ inequality only implies that the penalized log-likelihood increases at least as much as its expected complete-data version. Therefore, one can expect complex strategies to outperform simple ones (for instance, the hierarchical clustering strategy should outperform the relative eigengap strategy, that it encompasses, as discussed in the end of Section 2) but one cannot really prove it. The empirical experiments of the next section aim to better understand the respective advantages of each strategy.
@@ 5 Experiments
In this section, we perform various density modeling and clustering experiments with our new MPSA models (MPSA-H for the hierarchical clustering strategy, MPSA-R for the relative eigengap strategy, MPSA-U for the bottom-up strategy and MPSA-D for the top-down strategy), and compare them to the more standard full Gaussian mixture model (GMM-F) and the spherical one (GMM-S). The code is publicly available on GitHub: https://github.com/tomszwagier/MPSA. Technical implementation details can be found in the appendix (sections B.4 and B.6).
@@
Remark 9 (Choice of $\alpha$).
The penalty factor is set to $\alpha=\ln(n)/2$ in all the experiments, which corresponds to a BIC-like penalized log-likelihood. The choice of $\alpha$ certainly has an influence on the results—we can naturally expect the number of parameters of the selected model to decrease with $\alpha$. We conduct a sensitivity analysis in Section B.1, showing notably that one recovers the full GMM in the small $\alpha$ limit and the spherical GMM in the large $\alpha$ limit, and that the BIC tends to select (slightly) underparameterized models while the AIC tends to select (slightly) overparameterized models.
Remark 10 (Choice of $C$).
The number of clusters is set to the true one (if known) in the following experiments. Indeed, this paper focuses on the parsimony induced by equalizing close covariance eigenvalues rather than reducing the number of mixture components. However, for completeness, we discuss in Section B.2 how $C$ can be inferred asymptotically via a model selection framework or alternative methodologies relying on Dirichlet-type priors on the mixture proportions.
@@ 5.1 CPEM in action for 2D data
This experiment aims at understanding how the mixture parameters as well as the types evolve during the learning (Algorithm 2) with a simple two-dimensional example. For that, we take $n=1000$, $p=2$, $C=3$, $(\gamma_{1},\gamma_{2},\gamma_{3})=((1,1),(2),(2))$} (i.e. one anisotropic and two isotropic Gaussians), $(\pi_{1},\pi_{2},\pi_{3})=(0.4,0.3,0.3)$, $\mu_{c}\sim\mathcal{U}\left(\left[-8,8\right]^{p}\right)$, $(\lambda_{11},\lambda_{21},\lambda_{31})=(1,.5,.1)$, and $\lambda_{cp}/\lambda_{c1}=0.01\,\forall c$, and we fit an MPSA model with the hierarchical clustering strategy. We display in Figure 1 the scatter plot of the data and the learned mixture model, as well as the evolution of the penalized log-likelihood and the number of parameters.
@@ 5.2 Model selection strategies
The aim of this experiment is to study the learning algorithm 2 for different model selection strategies (MPSA-H/R/U/D) in higher dimension. To that extent, we plot in Figure 2 the penalized log-likelihood curves and the evolution of the number of parameters over the CPEM iterations, on a synthetic dataset with $n=1000$, $p=20$, $C=3$, $(\gamma_{1},\gamma_{2},\gamma_{3})=((3,5,12),(8,12),(20,))$}, $(\pi_{1},\pi_{2},\pi_{3})=(0.4,0.3,0.3)$, $\mu_{c}=0$, $(\lambda_{11},\lambda_{21},\lambda_{31})=(1,.5,.1)$, and $\lambda_{cp}/\lambda_{c1}=0.01\,\forall c$, with 10 independent repetitions.
Figure 2: Evolution of the penalized log-likelihood and the number of parameters over the iterations of the CPEM algorithm for the four model selection strategies (MPSA-H/R/U/D). Each transparent curve refers to one repetition while the thick opaque curve represents the average over the 10 independent repetitions.
We see that the penalized log-likelihoods are once again monotonous, as expected by theory (cf. Theorem 4).
Let us now comment the results based on our intuitions. All penalized log-likelihood curves seem to converge to the same optimum, after approximately 5 to 20 iterations. MPSA-D takes a much larger number of iterations to converge than the other strategies. However, the learning curves (penalized log-likelihood and number of parameters) are quite regular and with small variance, compared to the other strategies. We guess that in higher dimensions, MPSA-D may badly behave at first iterations due to the initialization with full covariance matrices, and it can moreover require a large number of iterations to converge, since it can only decrease the model complexity one by one. Therefore, smarter initialization heuristics could be designed, but this is left to future research. MPSA-U also yields quite regular learning curves, although we can observe small oscillations of the number of parameters near the optimum (cf. Remark 7 in Section 4.2). We moreover guess that MPSA-U may get stuck in local optima due to the simplicity of the models at first iterations [verbockhaven_growing_2024]. MPSA-H and MPSA-R are more flexible in terms of model complexity—which may make them attain the optimum faster as we can see in Figure 2. However, we guess that their flexibility may also cause higher variance in performance and parsimony.
5.3 Penalized log-likelihood benchmark
@@
[…]
Model MPSA $10$ MPSA $100$ Full $10$ Full $100$ Skew $100$ Ionosphere
@@
MPSA-R $\mathbf{-0.65\pm 0.06}$ $\mathbf{46.00\pm 0.22}$ $\mathbf{-07.59\pm 0.06}$ $100.02\pm 0.28$ $099.09\pm 0.72$ $\mathbf{+04.76\pm 1.67}$
@@
In the first setting (MPSA $10$), all our MPSA models significantly outperform the classical GMMs. Since $p=10$ and $n=1000$, one probably has enough samples to estimate the full (anisotropic) covariance parameters, and the additional parsimony of our MPSA models yields a significantly higher penalized log-likelihood. In the second setting (MPSA $100$), all our MPSA models again significantly outperform the classical GMMs. Since $p=100$ and $n=1000$, one probably does not have anymore enough samples to estimate the full covariance parameters, therefore the full GMM begins to severely drop in performance, while the spherical model is not flexible enough to model the data distribution due to its isotropic nature. In the third setting (Full $10$), since the true model is full and we have enough samples, the full GMM is the best estimator, and our MPSA models are on par with it, while the spherical GMM is below. In the fourth setting (Full $100$), the true model is still full but we don’t have enough samples anymore, so that the full GMM gets worse and the spherical GMM becomes better. Except for MPSA-R which is on par with GMM-S, our other MPSA models are significantly better. The same kind of observation goes for the fifth setting (Skew $100$). In the sixth setting (Ionosphere), our MPSA models significantly outperform the other GMMs.
@@
The previous experiments used the penalized log-likelihood (23) as a metric to evaluate the quality of the different GMMs. Beyond checking the monotonicity of the penalized log-likelihood over the CPEM iterations, these experiments showed that the MPSA models yield significantly higher penalized log-likelihoods than the classical GMMs—an empirical behavior that cannot be deduced from Theorem 4 alone. We believe that the penalized log-likelihood—as a tradeoff between goodness-of-fit and number of parameters—is a rather good indicator of the generalization properties of the learned models on unseen data, enjoying moreover statistical guarantees such as asymptotic consistency [keribin_consistent_2000, bai_consistency_2018]. That being said, it is clear that other interesting questions emanate from the study of the MPSA models, notably the one of density estimation: how well the MPSA models estimate the true underlying probability density. This is the purpose of the next subsection.
5.4 Density estimation benchmark
[…]
Figure 3: Density estimation with several GMMs. Several metrics are reported: the covariance errors, the type errors, the number of parameters and the (averaged) log-likelihood on an independent test set.
The component indexes $c$ of the true and estimated mixture densities are matched by minimizing the Euclidean distance between the centers; we check a posteriori that each estimated component is matched to one unique true component—i.e., the map corresponds to a permutation. Each metric is averaged over 10 independent experiments.
To summarize the results, everything behaves as expected: the estimation errors decrease with the number of samples and tend to zero. Interestingly, the covariance estimation errors are significantly smaller for the MPSA models than for the spherical and full GMMs—even for $n=1000$. This can be explained by looking more closely at the covariance eigenvalue profiles, which is what we do in Figure 4.
Figure 4: Covariance eigenvalue estimation with several GMMs for $n=1000$ (top) and $n=200$ (bottom). The three mixture components have, respectively, eigenvalue multiplicities $(15)$ (left), $(5,10)$ (middle) and $(5,5,5)$ (right)—plotted in black. The four MPSA strategies cannot be visually distinguished on the first five plots, as they yield relatively similar results. On the sixth plot, MPSA-H/U/D can be distinguished, while MPSA-R cannot be distinguished from GMM-S.
We see that the sample eigenvalues (represented by the purple curve) are highly dispersed around the population eigenvalues (represented by the black curve). This eigenvalue spreading phenomenon is well known in high dimensions (see, for instance, [johnstone_pca_2018]). In contrast, the eigenvalues estimated by the spherical GMM have low variance, but high bias for the second and third mixture component. The MPSA models not only achieve a significantly better estimation of the population eigenvalues, but also well capture the block structure, i.e., the eigenvalue multiplicities, which is not trivial. The eigenvalue plots of Figure 4 therefore well explain the results from Figure 3 in terms of covariance and type estimation.
We can see moreover that the number of parameters increases with the number of samples and seems to converge to the accurate one, testifying of the BIC’s asymptotic consistency. The spherical and full GMMs—respectively the most and least parsimonious GMMs belonging to the MPSA family—naturally bound the MPSA models on the plot of the number of parameters.
Finally, the test log-likelihood curves show that although the full GMM fails at estimating the covariance eigenvalues, it achieves similar density estimation results as the MPSA models, especially with respect to the spherical GMM.
Let us note that the estimation errors for the other mixture parameters—the proportions $\sum_{c=1}^{C}(\hat{\pi}_{c}-\pi_{c})^{2}$ and the means $\sum_{c=1}^{C}\|\hat{\mu}_{c}-\mu_{c}\|_{2}^{2}$—were also evaluated: the errors also tend to 0 for all models. We did not display them since the differences between the GMMs were not significant. Indeed, since the clusters are well separated, the proportions and means are well estimated by all models, independently of their covariance structure.
The differences between the model selection strategies (H/R/U/D) are not stringent in terms of estimation, at least in the setting of the experiment and for $n$ large enough, where the parameters are well estimated for all strategies. The MPSA-R strategy seems overall slightly worse than the other strategies, as we can particularly see on the tasks of covariance, type and eigenvalue estimation.